# State-observation augmented diffusion model for nonlinear assimilation with unknown dynamics

### v2

@@ 5 Conclusion and future work
Further improvements and extensions of our approach are also necessary. First, analyzing convergence rates and conditions in the presence of training errors would be beneficial. Robustness and stability are particularly important in real-world applications, and broader testing on both idealized and real-world models is needed to fully assess performance. Besides, adapting our SOAD approach to more complex tasks beyond data assimilation could also be an exciting avenue for future research.
@@ Acknowledgments
During the preparation of this work the authors used ChatGPT in order to polish the written text for spelling and grammar. After using this service, the authors reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
@@ Appendix A Configurations of the QG model
    Amplitude of the spectral spherical filter with parameter $23.6$ to obtain the small-scale dissipation;
@@
    Gravity: $g=9.81$;
@@
    Gradient of Coriolis parameter: $\beta=1.5\times 10^{-11}$;
@@
    Layer thickness: $H_{1}=500$, $H_{2}=2000$;
@@
    Upper/Lower layer flow: $U_{1}=0.025$, $U_{2}=0$.
Readers may refer to pyqg [^5] https://pyqg.readthedocs.io/en/latest/equations/notation_twolayer_model.html for the detailed numerical scheme.
@@ Appendix B Network architecture
The denoising network $\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)$ is implemented via a U-Net architecture proposed in [46], and we reproduce it in fig. 14 for completeness, where we have attached the channel numbers for each module.

### v3

@@ 5 Conclusion and future work
Further improvements and extensions of our approach are also necessary. First, analyzing convergence rates and conditions in the presence of training errors would be beneficial. Robustness and stability are particularly important in real-world applications, and broader testing on both idealized and real-world models is needed to fully assess performance. While our current framework assumes additive observational noise, extending it to handle non-additive or more complex noise structures would be a valuable direction for future work, as such cases may arise in certain practical scenarios. Besides, adapting our SOAD approach to more complex tasks beyond data assimilation could also be an exciting avenue for future research.
@@ Acknowledgments
During the preparation of this work the authors used ChatGPT in order to polish the written text for spelling and grammar. After using this service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@ Appendix A Configurations of the QG model
    Gravity: $g=9.81$;
@@
    Gradient of Coriolis parameter: $\beta=1.5\times 10^{-11}$;
@@
    Deformation radius: $r_{d}=1.5\times 10^{4}$;
@@
    Upper/Lower layer flow: $U_{1}=0.025$, $U_{2}=0$;
@@
    $F_{1}=r_{d}^{-2}/(1+\delta)$ and $F_{2}=\delta F_{1}$, where the layer thickness ratio $\delta=H_{1}/H_{2}=0.25$.
To obtain the small-scale dissipation “ssd” mentioned in eq. 45, a highly-selective exponential filter
${E_{f}=\begin{cases}\exp\left[-23.6(\kappa^{\star}-\kappa_{c})^{4}\right],&\kappa\geq\kappa_{c}\\ 1,&\kappa<\kappa_{c}\end{cases}}$
is employed as a multipler on the spectral domain. The term $\kappa^{\star}$ stands for the non-dimensional wavenumber, and the cutoff $\kappa_{c}$ is set as 65% of the Nyquist scale $\kappa_{ny}^{\star}=\pi$. Readers may refer to pyqg [^5] https://pyqg.readthedocs.io/en/latest/equations/notation_twolayer_model.html for the detailed numerical scheme.
@@ Appendix B Network architecture
The denoising network $\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)$ is implemented using a U-Net architecture proposed in [48], and we reproduce it in fig. 15 for completeness, where we have attached the channel numbers for each module.
@@
Appendix C Various observational noise distributions
Let $p(x)$ denote the probabilistic density function. To ensure all the observational noise distributions share the same mean and variance of our Gaussian assumptions $\mathcal{N}\left(0,(\sigma^{\mathrm{o}})^{2}\right)$ defined in eq. 51, we employ the following settings.
  • 1.
    Laplace:
    $p_{\mathrm{Laplace}}(x)=\frac{1}{2b}\exp\left(-\frac{|x-\mu|}{b}\right),$
    where $\mu=0$, $b=\sigma^{\mathrm{o}}/\sqrt{2}$;
  • 2.
    Uniform:
    $p_{\mathrm{Uniform}}(x)=\begin{cases}\frac{1}{2d},&|x|<d,\\ 0,&|x|\geq d,\end{cases}$
    where $d=\sqrt{3}\sigma^{\mathrm{o}}$;
  • 3.
    LogNormal: Let $Z\sim\mathcal{N}(\mu,s^{2})$, then the distribution of $\exp Z$ is defined to follow the LogNormal distribution with parameter $\mu$ and $s$. To make the random variable unbiased, we introduce an additional shift term $c$, which means we use $(\exp Z+c)$ to generate our observational noise. For any fixed $s$, we define
    $c=-\sigma^{\mathrm{o}}\left(\exp(s^{2})-1\right)^{-1/2},\quad\mu=-\frac{s^{2}}{2}+\log(-c).$
    The value of $s$ is appended to the distribution name in the main text.